OpenStudy (anonymous):

will medal !!!!!!!!!!!!!1 Which phrase could be represented by this expression? k/5 + 8 A. eight more than five times a number B. one-fifth of the sum of a number and eight C. eight more than a fifth of a number D. five times a number plus eight @Chibi_Robo3 @MARC_

OpenStudy (anonymous):

expression incomplete i guess?

OpenStudy (misty1212):

copy and paste fail

OpenStudy (misty1212):

\[5x+8\] maybe ?

OpenStudy (misty1212):

\[\frac{x}{5}+8\]?

OpenStudy (anonymous):

there

OpenStudy (alexandervonhumboldt2):

C

OpenStudy (alexandervonhumboldt2):

8+1/5k=k/5+8

OpenStudy (misty1212):

eight more than one fifth (of a number, not of whiskey)

OpenStudy (anonymous):

thankyou @misty1212

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (misty1212):

you are welcome dear